# Adversarial Agents: Black-Box Evasion Attacks with Reinforcement Learning

### v1

@@ 4.4 RL Attack Generalization and Comparison
Generalization Here we evaluate whether the attack generalizes to unseen inputs, i.e., samples not used for training. In Section 4.1, we construct disjoint sets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ for training and testing, respectively, with $5000$ images of support each. After training the attacks, we randomly sample 100 inputs from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ and evaluate the trained policies to record the misclassification, query count, and $\ell_{2}$-norm distortion. We repeat this experiment for 10 trials.
@@
In Table 1, we record the Misclassification Rate (MR), Median Queries (MQ), and Average $\ell_{2}$-norm distortion (Avg. $\ell_{2}$) for each trained RL attack type against train and test datasets $\mathcal{D}$ and $\mathcal{D^{\prime}}$. We see that the adversarial example metrics are very similar when comparing training to testing dataset performance. We perform a two-tailed Mann Whitney U-test to determine if the performance on the test dataset (across each metric) is out-of-distribution of the training dataset. We record the p-values of the hypothesis tests on each metric for each attack and observe that p-values range from 0.20 to 0.92. This means the performance on the test dataset $\mathcal{D}^{\prime}$ is well within the distribution of the train dataset $\mathcal{D}$ performance. Thus, the policy learned by the agent generalizes to unseen inputs.
Dataset RL Max Loss RL Min Norm
MR MQ Avg. $\ell_{2}$ MR MQ Avg. $\ell_{2}$
$(H_{0})$ Train Dataset 0.75$\pm$0.07 549 0.37$\pm$0.02 0.87$\pm$0.05 594 0.41$\pm$0.03
$(H_{1})$ Test Dataset 0.75$\pm$0.06 515 0.36$\pm$0.01 0.85$\pm$0.04 647 0.42$\pm$0.03
p-value ()($\neq$)( ≠ ) 0.92 0.32 0.24 0.20 0.33 0.24
Table 1: Comparison of RL attack types on train and test datasets showing misclassification rate (MR), median queries (MQ), and average $\ell_{2}$-norm distortion (Avg. $\ell_{2}$). The p-value results from the Mann-Whitney U-test that the distribution on the test dataset ($H_{1}$) is not equal ($\neq$) to that of the train dataset ($H_{0}$) for MR, MQ, and Avg. $\ell_{2}$.
Comparison Given the RL attacks’ ability to improve adversarial example performance in black-box settings, we compare the stream of inputs adversarial setting against SquareAttack. Like the MDP setup of the RL agent, we randomly sample an input from dataset $\mathcal{D}$ and apply SquareAttack to record the misclassification, query count, and $\ell_{2}$-norm distortion. For a fair comparison, we denote this as a single episode analogous to the RL attacks. We plot the rolling average misclassification rate, median queries, and $\ell_{2}$ norm distortion for two variations of SquareAttack: $\epsilon\in\{0.4,0.45\}$ against RL Max Loss ($\epsilon=0.5$) and RL Min Norm ($c=1.0$) across each metric in Figure 6.
(a) Misclassification Rate
(b) Median Queries
(c) Average $\ell_{2}$-norm Distortion
Figure 6: SquareAttack ($\epsilon=0.45$), SquareAttack ($\epsilon=0.4$), RL Max Loss, and RL Min Norm attack comparison with respect to episodes for 10 trials per attack with a 95% confidence interval.
On average, the SquareAttack misclassification rate, median queries, and average $\ell_{2}$-norm distortion is (0.691, 784, 0.425) throughout the entire stream of episodes, respectively. For comparison, the RL attacks in the first 100 episodes average (0.656, 1279, 0.361) and in the last 100 episodes average (0.822, 831, 0.411). This shows that by learning adversarial examples through RL, the adversarial agent succeeds 13.1% more after $5000$ consecutive attacks than SquareAttack, shown in 6(a). Further, SquareAttacks constant performance throughout the stream of episodes is an artifact of why an adversary is likely to use an RL approach in a sustained attack setting.
@@ 5 Discussion
MDP Improvements The RL attacks are only as good as the action space and reward function that allow the agent to change the state of the input with more reward. The current formulation selects $N$ features of an input with each selected feature getting $\delta_{i},i\in\{1,...,N\}$ with bounded distortion $|\delta_{i}|\leq\theta$ where $\theta\in(0,1)$ for the image domain. Distortion strategies come in many forms in AML, and alternative representations of the action space such as latent-space distortions could enhance the efficiency and attack success in RL. Further, including curriculum learning or adaptive rewards based on attack success could accelerate training. Future work should explore the combinations of different MDP components to develop more practical, real-world adversarial testing.
Cross-Domain applicability The study in this paper focuses on image classifiers, but the framework and analysis is not inherently tied to it. Images provide a natural domain to evaluate new attacks due to well-defined distortion metrics like the $\ell_{2}$-norm and rudimentary constraints the require features to be within $[0,255]$. Nevertheless, these strategies can be readily applied to other classification tasks such as malware and network intrusion detection. Investigating these domains not only studies the domain generalizability of the RL attacks proposed in this paper but also expands their impact on securing a range of machine learning applications.
Relevant connections to RL use in secruity RL has been explored in various security domains, including malware detection (Etter et al., 2023), network intrusion detection (Apruzzese et al., 2020), and image classification (Akers & Barton, 2024; Tsingenopoulos et al., 2019). These works rely on sparse reward functions and discrete action spaces, limiting the adaptability of learned policies. Moreover, these studies primarily evaluate RL-based attacks on a per-sample basis, optimizing perturbations for a fixed input without explicitly considering generalization across multiple examples.
@@ 6 Conclusion
This paper explores the capabilities of RL in black-box AML. We developed a novel MDP framework consisting of two agents, RL Max Loss and RL Min Norm, that emulate two classes of AML algorithms. The MDP setup enables the adversary to use RL to learn attack strategies that generate adversarial examples more successfully and with efficiency. Indeed, we find that: (1) the agents learn a policy that improves the misclassification rate by 19.4% and (2) the adversarial examples generated by the agents require 53.2% fewer queries over training on the CIFAR-10 dataset. Further, we analyze the sensitivity of reward hyperparameters $\epsilon$ and $c$ for RL Max Loss and RL Min Norm to show their consistency with state-of-the-art AML algorithms. Last, we demonstrate the trained policy’s ability to generalize crafting adversarial examples to inputs outside of the training dataset and show that in a comparison with a state-of-the-art black-box attack, the RL attacks become 13.1% more successful at generating adversarial examples after $5000$ episodes of training. These findings suggest that adversarial examples can be learned through RL, and that with the improvement of RL as a field results in stronger adversaries.
Broader Impacts
Using RL to craft adversarial examples has significant implications beyond evaluating model robustness. As AI becomes increasingly integrated with healthcare, finance, and cybersecurity, it is imperative that we understand their vulnerabilities and how an adversary can abuse them. RL attacks provide a dynamic way to evaluate models that static methods like traditional AML methods may miss. However, this raises ethical considerations regarding the misuse of these techniques. We emphasize the need for responsible disclosure and robust defense strategies. Future research should balance the offensive advancements with defensive improvements to ensure RL based attacks contributes to better AI security.
Appendix
Appendix A Model Architecture Details
PPO Parameter Value
Feature Extractor ResNet18
Policy Network Linear(128,64,$2N$)
Value Network Linear(128,64,1)
Activation ReLU
Learning Rate (LR) 2.5e-3×$\times\alpha$
$\gamma$ 0.99
GAE $\lambda$ 0.95
Clip Range 0.1×$\times\alpha$
Value Function Coef 0.5
Entropy Coef 0.01
Horizon $T$ 4096
Environments 32
@@ Appendix A Model Architecture Details
Batch Size 64
Epochs 10
Victim Parameter Value
Architecture ResNet50
Loss Function CrossEntropyLoss
@@
Weight Decay 5e-4
LR Scheduler Cosine Annealing
Batch Size 256
Epochs 50
Test Dataset Accuracy 95.67%
Table 2: PPO algorithm and victim model hyperparameters. $\alpha$ is linearly annealed from 1 to 0 over the course of training in PPO.
Acknowledgments
The authors thank Rahul Chatterjee, Yohan Beugin, and Kunyang Li for their helpful comments in previous iterations of this work.
Funding Acknowledgment. This material is based upon work supported by the National Science Foundation under Grant No. CNS-2343611. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Supplementary Materials
The following content was not necessarily subject to peer review.
Appendix B Action Hyperparameters
Figure 7: Misclassification Rate, Median Queries, and Average $\ell_{2}$-norm Distortion on adversarial examples post-training for different $(N,\theta)$ configurations with RL Min Norm and RL Max Loss attacks.
Appendix C RL Generated Adversarial Examples
Figure 8: CIFAR-10 adversarial examples generated by black-box RL attacks. Each image contains the confidence on the original class and confidence on the incorrect class.

### v2

@@ 4.4 Utility Analysis
Generalization. Here we evaluate whether the attack generalizes to unseen inputs, i.e., samples not used for training. Our experimental setup constructs disjoint sets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ for training and testing, respectively. After training the attacks, we evaluate the trained policies on each sample from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ on attack success, query count, and $\ell_{2}$-norm distortion. In Table 2, we record the ASR, AQ, and $\ell_{2}$ distortion for each trained RL attack type against train and test datasets $\mathcal{D}$ and $\mathcal{D^{\prime}}$ from CIFAR-10 and SVHN.
@@
We see that adversarial sample metrics are similar in distribution when comparing performance from training data $\mathcal{D}$ to testing data $\mathcal{D}^{\prime}$ in both CIFAR-10 and SVHN tasks across all victim models. This means the performance on the test dataset $\mathcal{D}^{\prime}$ is well within the distribution of the train dataset $\mathcal{D}$ performance. Thus, the policy learned by the agent generalizes to well to unseen inputs.
Comparison. Given the RL attacks’ ability to improve adversarial sample performance in black-box settings and generalize to new data, we compare adversarial agent generalization against known black-box attack baselines.We selected Square [andriushchenko_square_2020], HopSkipJumpAttack (HSJA) [chen_hopskipjumpattack_2020], and the Bandits [ilyas_prior_2019] attack, as these are prominent, state-of-the-art methods that represent different strategies for solving the black-box optimization problem (e.g., random search, decision-based, and prior-guided). To this end, we measure their ability to successfully craft adversarial samples on test dataset $\mathcal{D}^{\prime}$ from CIFAR-10 and SVHN by reporting ASR, AQ, and average $\ell_{2}$ distortion in the last three rows of Table 2. Indeed, we see that the ASR and AQ of both RL Max Loss and RL Min Norm on test dataset $\mathcal{D}^{\prime}$ outperform the most competitive baseline in most tasks with as much or fewer $\ell_{2}$ distortion. Notably, achieving up to 17% more attack success with 31% fewer queries in the most extreme case (VGG-16 on SVHN). Adversarial agent performance on new, unseen data is an artifact of why an adversary is likely to use an RL approach in a sustained attack setting where experience improves general attack performance.
@@ 5 Discussion
Sample Transferability. A key area of black-box AML research is model transferability: the degree to which adversarial samples on one model are adversarial in another. In this work, we investigate an agentic approach to query-based black-box attacks, focused on whether a RL policy (the agent) can learn a robust strategy for generating attacks. Our results demonstrate that this is effective and generalizable to traditional image classification benchmarks and state-of-the-art image classification models. These findings suggest that adversarial agents can learn attack algorithms through RL. This RL framework recasts attack transferability as a more promising opportunity with policy adaptation. The extensive training of the RL policy on an initial victim model can be viewed as an effective pre-training step. We hypothesize that this pre-trained RL policy could then be adapted (or fine-tuned) using the same RL framework on a new victim model with greater attack success and query efficiency than a policy trained from scratch. This approach remains a logical and compelling direction for future work.
MDP improvements. The RL attacks are only as good as the action space and reward function that allow the agent to change the state of the input with more reward. The current formulation selects $N$ features of an input with each selected feature getting $\delta_{i},i\in\{1,...,N\}$ with bounded distortion $|\delta_{i}|\leq\theta$ where $\theta\in(0,1)$ for the image domain. Distortion strategies come in many forms in AML, and alternative representations of the action space such as latent-space distortions could enhance the efficiency and attack success in RL. Further, including curriculum learning or adaptive rewards based on attack success could accelerate training. Future work should explore the combinations of different MDP components to develop more practical, real-world adversarial testing.
Cross-domain applicability. The study in this paper focuses on image classifiers, but the framework and analysis is not inherently tied to it. Images provide a natural domain to evaluate new attacks due to well-defined distortion metrics like the $\ell_{2}$-norm and rudimentary constraints that require features to be within $[0,255]$. Nevertheless, these strategies can be readily applied to other classification tasks such as malware and network intrusion detection. Investigating these domains not only studies the domain generalizability of the RL attacks proposed in this paper but also expands their impact on securing a range of machine learning applications.
@@ 6 Conclusion
This paper explores the capabilities of RL in black-box AML. We developed a novel MDP framework consisting of two agents, RL Max Loss and RL Min Norm, that emulate two classes of AML algorithms. The MDP setup enables the adversary to use RL to learn attack strategies that generate adversarial samples more successfully and with efficiency. Indeed, we find that: (1) the agents learn a policy that improves the attack success rate by up to 13.2% and (2) the adversarial samples generated by the agents require up to 16.9% fewer queries over training on benchmark image classification datasets. Further, we analyze the sensitivity of reward hyperparameters $\epsilon$ and $c$ for RL Max Loss and RL Min Norm to show their consistency with state-of-the-art AML algorithms. Last, we demonstrate the trained policy’s ability to generalize crafting adversarial samples to inputs outside of the training dataset and show that in a comparison with state-of-the-art black-box attacks, the RL attacks become 17% more successful at generating adversarial samples on unseen inputs post-training. These findings suggest that better black-box evasion attack strategies can be learned through RL, and that with the improvement of RL as a field results in stronger adversaries.
Supplementary Material
This supplementary material provides additional details on the experimental setup, including the selection of action hyperparameters and qualitative examples of generated adversarial samples.
Action Hyperparameters
Figure 6: Misclassification Rate, Median Queries, and Average $\ell_{2}$-norm Distortion on adversarial examples post-training for different $(N,\theta)$ configurations with RL Min Norm and RL Max Loss attacks on the CIFAR-10 dataset.
In Section 3.1, we define the action $a_{t}$ as a set of $N$ feature-perturbation pairs, $\{(i_{1},\delta_{1}),...,(i_{N},\delta_{N})\}$, where each perturbation $\delta_{j}$ has a maximum magnitude of $\theta$. The choice of $N$ (the number of features to perturb) and $\theta$ (the maximum magnitude of that perturbation) balances the trade-off between attack success and the complexity of the action space.
To select appropriate values, we conducted preliminary experiments, the results of which are shown in Figure 6. This figure plots the misclassification rate (ASR), median queries, and average $\ell_{2}$-norm distortion for various $(N,\theta)$ configurations for both RL Max Loss and RL Min Norm attacks on CIFAR-10. Based on these results, we fixed $N=5$ and $\theta=0.05$ for all main experiments in the paper, as this configuration offered a good balance between attack success and action complexity.
Hyperparameters
Parameter Value
PPO Agent (EfficientNet)
Policy/Value Arch. Linear(128, 64)
@@
Learning Rate (LR) 2.5e-3
Discount Factor ($\gamma$) 0.99
GAE Lambda ($\lambda$) 0.95
Clip Range 0.1
Victim Models (ResNet50, VGG16)
@@
Learning Rate (LR) 0.001
LR Scheduler ReduceLROnPlateau
Weight Decay 1e-4
Victim Model (ViT_B_16)
Optimizer AdamW
Learning Rate (LR) 0.001
LR Scheduler ReduceLROnPlateau
Weight Decay 1e-4
Table 3: Key hyperparameters for the PPO agent and the victim models. For the PPO agent, the learning rate and clip range are linearly annealed from their initial values to 0 over the course of training.
Table 3 details the key hyperparameters used for training both the PPO agent and the victim models. The PPO agent’s policy and value functions use an EfficientNet feature extractor. The victim models were fine-tuned from ImageNet-1K pre-trained weights.
RL Generated Adversarial Samples
Figure 7: CIFAR-10 adversarial samples generated by black-box RL attacks. Each image contains the confidence on the original class and confidence on the incorrect class.
To provide a qualitative sense of the attacks, Figure 7 visualizes several adversarial samples generated by our trained RL agents on the CIFAR-10 dataset. Each example shows the original image, the resulting adversarial image, and the imperceptible perturbation (magnified for visibility). The labels demonstrate the agent’s success: the victim model’s confidence is shifted from the high-confidence original class to a high-confidence incorrect class. This aligns with the overall framework described in Figure 1, where the RL adversary iteratively queries the victim model to produce an adversarial example.